# Partial identification and unmeasured confounding with multiple treatments and multiple outcomes

### v2

@@ Abstract
In this work, we develop a framework for partial identification of causal effects in settings with multiple treatments and multiple outcomes. We highlight several advantages of jointly analyzing causal effects across multiple estimands under a “factor confounding assumption” where residual dependence amongst treatments and outcomes is assumed to be driven by unmeasured confounding. In this setting, we show that joint partial identification regions for multiple estimands can be more informative than considering partial identification for individual estimands one at a time. We additionally show how assumptions related to the strength of confounding or the magnitude of plausible effect sizes for any one estimand can reduce the partial identification regions for other estimands. As a special case of this result, we explore how negative control assumptions reduce partial identification regions and discuss conditions under which point identification can be obtained. We develop novel computational approaches to finding partial identification regions under a variety of these assumptions. Lastly, we demonstrate our approach in an analysis of the causal effects of multiple air pollutants on several health outcomes in the United States using claims data from Medicare, where we find that some exposures have effects that are robust to the presence of unmeasured confounding.
@@ 1 Introduction
Causal inference in observational studies is a crucially important methodological framework to help provide policy-relevant evidence in situations where experimentation or randomization is not possible. In these situations, untestable assumptions must be made, and researchers should, to the best of their ability, evaluate how robust their conclusions are to violations of those assumptions. While there are many different assumptions that can be made in order to obtain causal identification in different scenarios, one of the most commonly employed is that of no unmeasured confounding, which states that the observed covariates contain all common causes of the treatment and outcome. In the presence of unmeasured confounding, causal effects are not identifiable, and consistent estimation of them is not possible without additional information or assumptions. In this work, we focus on approaches to partial identification of causal effects when there is unmeasured confounding, with a particular focus on settings with multiple treatments and multiple outcomes, which we show has many benefits for partial identification.
@@
The most common approach for assessing the unmeasured confounding assumption is sensitivity analysis, which dates back at least as far as Cornfield et al. (1959) who argued that there is a causal relationship between smoking and lung cancer. A central goal of sensitivity analysis is to provide bounds on causal estimands that account for potential biases stemming from unmeasured confounding. In simple cases such as with bounded outcomes, general bounds for causal effects can be estimated (Manski, 1990), though typically bounds are estimated based on sensitivity parameters governing the strength of association between the unmeasured confounders and either the treatment or outcome. One approach is to posit a sensitivity parameter for the relationship between the unmeasured variable and treatment, while assuming the worst-case scenario that the unmeasured confounder and outcome are nearly co-linear (Rosenbaum, 1987, 1988, 1991, 2002b, 2002a). Other approaches have two sensitivity parameters that govern the strength of association between the unmeasured variable and both the treatment and outcome. These parameters are intended to be interpretable so that one can reason potential values for them. This can be done by letting them be parameters of a regression model (Rosenbaum & Rubin, 1983; McCandless et al., 2007), parameters dictating relative risks (VanderWeele & Ding, 2017; Ding & VanderWeele, 2016), or partial $R^{2}$ values between the unmeasured variable and the treatment or outcome (Imbens, 2003; Small, 2007; Veitch & Zaveri, 2020; Cinelli & Hazlett, 2020; Freidling & Zhao, 2022; Chernozhukov et al., 2022).
@@
While much of the existing work focuses on a single treatment variable, recent work has examined settings with multiple treatments (Miao et al., 2022; Kong et al., 2022; Zheng et al., 2021, 2022). Under certain assumptions about the nature of confounding, partial $R^{2}$ values between the unmeasured confounders and treatment can be identified from the observed data. In some cases, such as sparsity in the effects of the treatment on the outcome, the causal effects are even identifiable in this setting (Miao et al., 2022). Another example is Kong et al. (2022), which shows identifiability with multiple treatments and a binary outcome under the assumption the treatments are independent given a common unmeasured confounder. Similar results have been shown in the multi-outcome setting (Zhou et al., 2020; Zheng et al., 2023). Recent work in Wu et al. (2023) has extended these ideas in a number of directions, in particular to the multiple treatment and multiple outcome setting, where causal identification can be obtained under certain conditional independence assumptions related to those seen in the proximal causal inference literature (Tchetgen Tchetgen et al., 2024). In general, the presence of multiple treatments or outcomes can provide substantial benefits. As mentioned above, one can obtain identification in these settings even with unmeasured confounding, albeit under fairly strong, restrictive assumptions. Under weaker assumptions, one can potentially obtain partial identification of treatment effects, where only a set of values of the causal effect is identified, which is the focus of this work.
@@
In this paper, we develop a general partial identification framework for scenarios with multiple treatments and multiple outcomes. Under a factor confounding assumption, we are able to identify worst-case bounds for causal effects without any additional sensitivity parameters. Further, we show that additional assumptions about the bias, effect size or strength of confounding on any set of estimands can decrease the worst-case bounds for other causal estimands, highlighting the benefit of joint inference over one-at-a-time analysis. Specifically, we illustrate two different classes of assumptions which we can be used reduce the joint partial identification region for all causal estimands: 1) assumptions about the magnitude of effect sizes (or bias), or 2) assumptions about the strength of confounding, e.g., by bounding the fraction of outcome variance and treatment variance explained by unmeasured confounders (given measured confounders). We provide theoretical results for these strategies and develop novel computational tools for characterizing partial identification regions. Our work is motivated by an application exploring the health effects of multivariate air pollution exposures in a nationwide study in the United States using Medicare claims data. We utilize the aforementioned approaches to study whether the adverse health effects of air pollution are robust to unmeasured confounding, and how the partial identification regions for causal effects of interest vary across different assumptions about the strength of confounding.
@@ 2 Notation and estimands
Throughout, we observe $(Y_{i},T_{i},X_{i})$ for $i=1,\dots,n$. Here, $T=(T_{1},\cdots,T_{k})^{\top}$ denotes a vector of $k$ treatments, $Y=(Y_{1},\cdots,Y_{q})^{\top}$ denotes a vector of $q$ outcomes, and $X=(X_{1},\cdots,X_{l})^{\top}$ denotes a vector of $l$ pre-treatment covariates, respectively. Under the potential outcome framework (Rubin, 1974; Splawa-Neyman et al., 1990), we let $Y(t)$ denote the potential outcome that would have been observed had the treatment $T$ been set to $t$. Our goal throughout will be to utilize the observed data to estimate the population average treatment effect (PATE), for a linear combination of outcomes $a^{\top}Y$, defined as
@@
for some $t_{1},t_{2}$. If unmeasured confounders $U$ were observed along with $X$, we could identify the causal effect from the observed data under the following assumptions:
Assumption 1 (SUTVA; Stable Unit Treatment Value Assumption).
There is no interference between units and there is only a single version of each assigned treatment.
Assumption 2 (Latent ignorability).
$Y(t)\perp\!\!\!\perp T\mid(U,X)$ for all $t$.
Assumption 3 (Latent positivity).
$f(t\mid U=u,X=x)>0$ for all possible values of $t$, $x$, and $u$. Here, $f(\cdot)$ represents the conditional density function of $T$ given $(X,U)$.
Assumption 1 ensures that $Y(t)$ is well-defined and that $Y=Y(t)$ if $T=t$, which links the observed data to potential outcomes. Assumption 2 states that $X$ and $U$ contain all common causes of $T$ and $Y$. Under this assumption, we can identify the causal effect of interest using $E[Y(t)]=E_{X,U}[E(Y|T=t,X,U)]$. Assumption 3, also referred to as overlap, assures that there is a positive probability that each unit receives any treatment value $t$ given both $X$ and $U$.
@@ 3.1 Factor models with multivariate treatments and outcomes
For ease of exposition, we do not include observed covariates $X$ in this section, though all results hold analogously when conditioning on $X$. Throughout, we consider the following latent variable models for the observed and unobserved data:
$U$ $=\epsilon_{U},~{}~{}~{}$ (1)
$T$ $=BU+\epsilon_{T},~{}~{}~{}$ (2)
$Y$ $=g(T)+\Gamma\Sigma_{u|t}^{-1/2}U+\epsilon_{Y},~{}~{}~{}$ (3)
where $B\in\mathbb{R}^{k\times m}$, $\Gamma\in\mathbb{R}^{q\times m}$, and $\Sigma_{u|t}:=\text{Cov}(U|T=t)$. We let $\epsilon_{U}\sim N_{m}(0,I_{m})$, $\epsilon_{T}\sim N_{k}(0,\Sigma_{t|u})$, and $\epsilon_{Y}\sim N_{q}(0,\Sigma_{y|t,u})$. We assume that $\epsilon_{Y}\perp\!\!\!\perp T,U$. From (1)–(2), the conditional confounder distribution can be derived as
@@
Our interest is to estimate $E[Y(t)]=g(t)$, which we want to recover from the observed data. The observed outcome distribution, however, has mean and variance given by
$E(Y|T=t)$ $=\check{g}(t)=g(t)+\Gamma\Sigma_{u|t}^{-1/2}\mu_{u|t},$
$\text{Var}(Y|T=t)$ $=\Sigma_{y|t}=\Gamma\Gamma^{\top}+\Sigma_{y|t,u}.$ (5)
This shows that our estimand of interest and its corresponding bias when ignoring $U$ can be expressed as
$\text{PATE}_{a,t_{1},t_{2}}$ $:=E[a^{\top}Y(t_{1})-a^{\top}Y(t_{2})]$ (6)
$=a^{\top}g(t_{1})-a^{\top}g(t_{2})$ (7)
and
$=a^{\top}\Gamma\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t},$ (8)
where $\mu_{u|\Delta t}=\mu_{u|t_{1}}-\mu_{u|t_{2}}$.
3.2 Bounding the confounding bias
Theorem 1.
Assuming models (1)–(3), the confounding bias of $\text{PATE}_{a,t_{1},t_{2}}$ can be bounded as
$\text{Bias}_{a,t_{1},t_{2}}^{2}\leq\|a^{\top}\Gamma\|_{2}^{2}\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|^{2}_{2},$
The proof of the theorem is given in Appendix A.1. $a^{\top}\Gamma$ can be viewed as the strength of association between the unmeasured variables $U$ and the outcome, while $\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$ is the scaled difference in unmeasured confounder means when treatment is shifted from $t_{1}$ to $t_{2}$ and represents the strength of association between the unmeasured confounders and treatment. This result implies that the true treatment effect lies in the interval
$a^{\top}[\check{g}(t_{1})-\check{g}(t_{2})]\pm\|a^{\top}\Gamma\|_{2}\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2},$ (9)
which is a partial identification region for the PATE and is centered at the naïvely estimated PATE. While the confounding bias itself is not identifiable, it can be shown that this bound is identified under an assumption on the unmeasured confounders, which we detail in the following assumption.
@@ 3.2 Bounding the confounding bias
The data is generated by (1)–(3) and $U$ represents potential confounders in the sense that they are possible causes of $T$ and $Y$ and are not themselves caused by $T$ or $Y$. Additionally, we assume the following conditions are satisfied.
  1. (C1)
@@
Given Assumptions 1-4, both $\|a^{\top}\Gamma\|_{2}^{2}$ and $\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|^{2}_{2}$ are identifiable and thus the bound on $\text{Bias}_{a,t_{1},t_{2}}$, $\|a^{\top}\Gamma\|_{2}^{2}\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|^{2}_{2}$, is also identifiable. Further, for any orthogonal matrix $\widetilde{\Gamma}$ satisfying $\widetilde{\Gamma}\widetilde{\Gamma}^{\top}=\Gamma\Gamma^{\top}$ and $\tilde{B}$, satisfying $\widetilde{B}\widetilde{B}^{\top}=BB^{\top}$ there exists an $m\times m$ orthonormal matrix, $R$, such that $\text{Bias}_{a,t_{1},t_{2}}=a^{\top}\widetilde{\Gamma}R\widetilde{\Sigma}_{u|t}^{-1/2}\widetilde{\mu}_{u|\Delta t}$, where $\widetilde{\mu}_{u|\Delta t}$ and $\widetilde{\Sigma}_{u|t}^{-1/2}$ are the matrices implied by the particular value of $\tilde{B}$. The equivalence class of matrices $\widetilde{\Gamma}$ and $\widetilde{B}$ are identifiable, so the only unidentified component of the bias is $R$.
@@
Note that Assumption 4 states that $U$ are in fact confounders. The factor models are capturing correlation in the residuals of the treatment and outcome and while we are assuming this comes from confounders, it could also come from predictors of either treatment or outcome, as well as mediators that occur post-treatment. The bias bound in Theorem 1 is still valid in such settings, but is likely to be conservative as we will over-estimate the extent to which unmeasured confounding could bias our estimates.
4 Partial identification with additional constraints

### v3

@@ Abstract
Estimating the health effects of multiple air pollutants is a crucial problem in public health, but one that is difficult due to unmeasured confounding bias. Motivated by this issue, we develop a framework for partial identification of causal effects in the presence of unmeasured confounding in settings with multiple treatments and multiple outcomes. Under a factor confounding assumption, we show that joint partial identification regions for multiple estimands can be more informative than considering partial identification for individual estimands one at a time. We show how assumptions related to the strength of confounding or magnitude of plausible effect sizes for one estimand can reduce the partial identification regions for other estimands. As a special case of this result, we explore how negative control assumptions reduce partial identification regions and discuss conditions under which point identification can be obtained. We develop novel computational approaches to finding partial identification regions under a variety of these assumptions. We then estimate the causal effect of PM2.5 components on a variety of public health outcomes in the United States Medicare cohort, where we find that, in particular, the detrimental effect of black carbon is robust to the potential presence of unmeasured confounding bias.
@@ 1 Introduction
Understanding the health effects of air pollution is a critically important problem in public health research. There is a vast literature that indicates substantial detrimental impacts of air pollution (Cohen et al.,, 2017; Burnett et al.,, 2018; Manisalidis et al.,, 2020). This has helped inform regulatory policy on how to reduce air pollution, and it is important to obtain accurate estimates of the impact of these policies as both the costs and benefits of these policies are in the billions of dollars (Portney,, 1990). Due to the importance of this problem and the need for policy-relevant estimates of the impacts of air pollution, there has been a push towards causal inference approaches in environmental epidemiology (Dominici and Zigler,, 2017; Carone et al.,, 2020; Sommer et al.,, 2021). Causal inference in these settings is difficult for a multitude of reasons, but arguably the biggest impediment to assessing the causal effect of environmental exposures is the potential for unmeasured confounding bias. In the presence of unmeasured confounding, causal effects are not identifiable, and consistent estimation of them is not possible without additional information or assumptions. In some settings, a natural experiment is available that helps to alleviate issues stemming from unmeasured confounding (Rich,, 2017). Examples include pollution reductions due to the Olympic games in Atlanta or Beijing (Huang et al.,, 2015) or the decommissioning of power plants leading to large reductions in pollution for nearby areas (Luechinger,, 2014). Other work has looked to leverage recent advancements in double negative control designs to estimate causal effects in a manner that is robust to unmeasured confounding (Schwartz et al.,, 2023). Generally, however, natural experiments or other designs that alleviate concerns about unmeasured confounding are not always available. Given the importance of this problem in air pollution research and its ubiquity in the broader causal inference framework, we develop methods for partial identification of causal effects under unmeasured confounding. We pay particular attention to settings with multiple treatments and outcomes, which we show offer significant advantages for partial identification.
@@
Historically, the most common approach for assessing the unmeasured confounding assumption is sensitivity analysis, which dates back at least as far as Cornfield et al., (1959) who argued that there is a causal relationship between smoking and lung cancer. A central goal of sensitivity analysis is to provide bounds on causal estimands that account for potential biases stemming from unmeasured confounding. In simple cases such as with bounded outcomes, general bounds for causal effects can be estimated (Manski,, 1990), though typically bounds are estimated based on sensitivity parameters governing the strength of association between the unmeasured confounders and either the treatment or outcome. One approach is to posit a sensitivity parameter for the relationship between the unmeasured variable and treatment, while assuming the worst-case scenario that the unmeasured confounder and outcome are nearly co-linear (Rosenbaum,, 1987, 1988, 1991; Rosenbaum, 2002b, ; Rosenbaum, 2002a, ). Other approaches have two sensitivity parameters that govern the strength of association between the unmeasured variable and both the treatment and outcome. These parameters are intended to be interpretable so that one can reason potential values for them. This can be done by letting them be parameters of a regression model (Rosenbaum and Rubin,, 1983; McCandless et al.,, 2007), parameters dictating relative risks (VanderWeele and Ding,, 2017; Ding and VanderWeele,, 2016), or partial $R^{2}$ values between the unmeasured variable and the treatment or outcome (Imbens,, 2003; Small,, 2007; Veitch and Zaveri,, 2020; Cinelli and Hazlett,, 2020; Freidling and Zhao,, 2022; Chernozhukov et al.,, 2022).
@@
While much of the existing work focuses on a single treatment variable, recent work has examined settings with multiple treatments (Wang et al.,, 2017; Miao et al.,, 2022; Kong et al.,, 2022; Zheng et al.,, 2021, 2022). Under certain assumptions about the nature of confounding, partial $R^{2}$ values between the unmeasured confounders and treatment can be identified from the observed data. In some cases, such as sparsity in the effects of the treatment on the outcome, the causal effects are even identifiable in this setting (Wang et al.,, 2017; Miao et al.,, 2022). Another example is Kong et al., (2022), which shows identifiability with multiple treatments and a binary outcome under the assumption the treatments are independent given a common unmeasured confounder. Similar results have been shown in the multi-outcome setting (Zhou et al.,, 2020; Zheng et al.,, 2023). Recent work in Wu et al., (2023) has extended these ideas in a number of directions, in particular to the multiple treatment and multiple outcome setting, where causal identification can be obtained under certain conditional independence assumptions related to those seen in the proximal causal inference literature (Kuroki and Pearl,, 2014; Tchetgen Tchetgen et al.,, 2024). In general, the presence of multiple treatments or outcomes can provide substantial benefits, which is particularly important for air pollution research as there are large numbers of environmental exposures of interest, and they are potentially associated with a variety of adverse health outcomes. As mentioned above, one can obtain identification in these settings even with unmeasured confounding, albeit under fairly strong, restrictive assumptions. Under weaker assumptions, one can potentially obtain partial identification of treatment effects, where only a set of values of the causal effect is identified, which is the focus of this work.
@@
In this paper, we develop a general partial identification framework for scenarios with multiple treatments and multiple outcomes to study whether the estimated health effects of air pollution are robust to the presence of unmeasured confounding bias. While motivated by settings in air pollution epidemiology, this work has broader implications for partial identification of causal effects with multiple treatments and outcomes. Under a factor confounding assumption, we are able to identify worst-case bounds for causal effects without any additional sensitivity parameters. Further, we show that additional assumptions about the bias, effect size or strength of confounding on any set of estimands can decrease the worst-case bounds for other causal estimands, highlighting the benefit of joint inference over one-at-a-time sensitivity analysis. Specifically, we illustrate two different classes of assumptions which we can be used reduce the joint partial identification region for all causal estimands: 1) assumptions about the magnitude of effect sizes (or bias), or 2) assumptions about the strength of confounding, e.g., by bounding the fraction of outcome variance and treatment variance explained by unmeasured confounders (given measured confounders). We provide theoretical results for these strategies and develop novel computational tools for characterizing partial identification regions. We then estimate the health effects of multivariate air pollution exposures in a nationwide study in the United States using Medicare claims data. We utilize the aforementioned approaches to study whether the adverse health effects of air pollution are robust to unmeasured confounding, and quantify how much the partial identification regions for causal effects of interest change under different assumptions about the strength of confounding.
2 Motivating study of the health effects of air pollution
Understanding the health effects of air pollution is a critically important problem in public health research. While there is a large literature showing adverse health effects of air pollution, nearly all of the research is observational, and therefore relies on certain untestable assumptions. A large number of these studies assume that the observed covariates contain all confounders of the exposure-outcome relationship. Due to the importance of this public health problem, and the impact these studies have on environmental regulatory policy, it is important to assess whether these findings are robust to unmeasured confounding bias. This manuscript works to develop methodology to estimate the causal effect of ozone and specific components of PM2.5 on a variety of health outcomes, while accounting for the possibility that these are subject to unmeasured confounding bias.
To study this question, we use claims data from Medicare Fee-for-Service enrollees over the age of 65 that were enrolled in the years 2000 to 2016 obtained from the Centers of Medicare and Medicaid Services. For each enrollee we have information on their age, sex, race, and whether they are Medicaid eligible, which functions as a proxy for low socioeconomic status, as well as their residential zip code. In addition to these individual level characteristics, we have a number of area level characteristics that are available through the United States Census Bureau and the Center for Disease Control’s Behavioral Risk Factor Surveillance System. These include covariates such as average body mass index, education level, population density, smoking rates, median household income and house value, and the percentage of owner occupied housing. We also adjust for meteorological variables, such as temperature and precipitation rates from the National Climatic Data Center.
In our analysis, we focus on six different air pollution exposures from two distinct publicly available sources: we obtain estimates of ammonium (NH4), nitrates (NO3), and sulfate (SO4) levels on a ($0.01^{\circ}\times 0.01^{\circ}$) grid from the Atmospheric Composition Analysis Group (Van Donkelaar et al.,, 2019) and we obtain estimates of elemental carbon (EC), organic carbon (OC), and ozone (O3) on a 1km by 1km daily grid from the Socioeconomic Data and Applications Center (Di et al.,, 2021; Requia et al.,, 2021). These exposures are aggregated to the zip code level so that they can be linked to the residential zip code obtained from Medicare, and are additionally aggregated to the yearly level as we are examining long term exposure to pollution. The outcomes of interest consist of zip code level hospitalization rates for anemia, COPD, lung cancer, stroke, and asthma, which are obtained based on ICD9/10 diagnosis codes. Additionally, we have access to negative control outcomes, which are hospitalization rates in the year prior to exposure. We focus on using hospitalization rates for COPD in the prior year as a negative control outcome. Our unit of analysis is a United States zip code, and our data consists of all zip codes in the contiguous United States. All covariates are aggregated up to the zip code level by taking their averages or proportions within each zip code. Similar data sources have been used previously to provide impactful, large-scale estimates of the health effects of PM2.5 (Di et al.,, 2017). Critically, these results have relied on the no unmeasured confounding assumption, and in this work we aim to clarify the extent to which these effects are robust to unmeasured confounding. This will provide environmental regulators with improved scientific evidence on the nature of the relationship between air pollution and public health.
@@ 3 Notation and estimands
Throughout, we observe $(Y_{i},T_{i},X_{i})$ for $i=1,\dots,n$. Here, $T=(T_{1},\cdots,T_{k})^{\top}$ denotes a vector of $k$ treatments, $Y=(Y_{1},\cdots,Y_{q})^{\top}$ denotes a vector of $q$ outcomes, and $X=(X_{1},\cdots,X_{l})^{\top}$ denotes a vector of $l$ pre-treatment covariates, respectively. Throughout, we use the terms exposures and treatments interchangeably when referring to $T$. Under the potential outcome framework (Rubin,, 1974; Splawa-Neyman et al.,, 1990), we let $Y(t)$ denote the potential outcome that would have been observed had the treatment $T$ been set to $t$. Our goal throughout will be to utilize the observed data to estimate the population average treatment effect (PATE), for a linear combination of outcomes $a^{\top}Y$, defined as
@@
for some $t_{1},t_{2}$. If unmeasured confounders $U$ were observed along with $X$, we could nonparametrically identify the causal effect from the observed data under the following assumption:
Assumption 1.
  1. 1.
    SUTVA (Stable Unit Treatment Value Assumption): There is no interference between units and there is only a single version of each assigned treatment.
  2. 2.
[…]
  3. 3.
    Latent positivity: $f(t\mid U=u,X=x)>0$ for all possible values of $t$, $x$, and $u$. Here, $f(\cdot)$ represents the conditional density function of $T$ given $(X,U)$.
The SUTVA assumption ensures that $Y(t)$ is well-defined and that $Y=Y(t)$ if $T=t$, which links the observed data to potential outcomes. The latent ignorability assumption states that $X$ and $U$ contain all common causes of $T$ and $Y$. The latent positivity assumption, also referred to as overlap, assures that there is a positive probability that each unit receives any treatment value $t$ given both $X$ and $U$. Under these assumptions, we can identify the causal effect of interest using $E[Y(t)]=E_{X,U}[E(Y|T=t,X,U)]$.
@@ 4.1 Factor models with multivariate treatments and outcomes
For ease of exposition, we do not include observed covariates $X$ in this section, though all results hold analogously when conditioning on $X$. Throughout, we consider the latent variable models that are described by the following assumptions.
Assumption 2 (Confounder homoskedasticity).
Let $E[U|T=t]:=\mu_{u|t}$ and $Var(U|T=t):=\Sigma_{u|t}$ be the $m$-dimensional mean and $m\times m$ covariance matrix of latent confounders given treatments. Assume that the confounders are homoskedastic in $t$, meaning that $\Sigma_{u|t}$ is invariant to the level of $t$.
Assumption 3 (Outcomes are linear in latent confounders).
The mean of the outcome is linear in the unmeasured confounders with mean
$E[Y|T,U]$ $=g(T)+\Gamma\Sigma_{u|t}^{-1/2}U.~{}~{}~{}$ (1)
with $\Gamma\in\mathbb{R}^{q\times m}$ and $q$-dimensional diagonal variance $Var(Y|T,U):=D$.
Note that if we had observed covariates $X$, they would be placed inside of the $g(\cdot)$ function along with $T$ thereby allowing for arbitrary relationships between the observed variables. We only require a specific structure for the effect of the unobserved confounders on the outcome. Assumptions 2 and 3 imply that that the observed outcome moments are
$E(Y|T=t)$ $=\check{g}(t)=g(t)+\Gamma\Sigma_{u|t}^{-1/2}\mu_{u|t},$ (2)
$\text{Var}(Y|T=t)$ $=\Gamma\Gamma^{\top}+D.$ (3)
Given that our estimand of interest is
$\text{PATE}_{a,t_{1},t_{2}}$ $:=E[a^{\top}Y(t_{1})-a^{\top}Y(t_{2})]$ (4)
$=a^{\top}g(t_{1})-a^{\top}g(t_{2}),$ (5)
Equation (2) shows that that the bias from ignoring unmeasured confounders, $U$ is
$=a^{\top}\Gamma\Sigma_{u|t}^{-1/2}(\mu_{u|t_{1}}-\mu_{u|t_{2}}).$ (6)
Theorem 1.
Given Assumptions 13, the confounding bias of $\text{PATE}_{a,t_{1},t_{2}}$ can be bounded as
$\text{Bias}_{a,t_{1},t_{2}}^{2}\leq\|a^{\top}\Gamma\|_{2}^{2}\|\Sigma_{u|t}^{-1/2}(\mu_{u|t_{1}}-\mu_{u|t_{2}})\|^{2}_{2},$
The proof of the theorem is given in Appendix A. $a^{\top}\Gamma$ can be viewed as the strength of association between the unmeasured variables $U$ and the outcome, while $\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$ is the scaled difference in unmeasured confounder means when treatment is shifted from $t_{1}$ to $t_{2}$ and represents the strength of association between the unmeasured confounders and treatment. This result implies that the true treatment effect lies in the interval
$a^{\top}[\check{g}(t_{1})-\check{g}(t_{2})]\pm\|a^{\top}\Gamma\|_{2}\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}.$ (7)
While the confounding bias itself is not identifiable, it can be shown that this bound is identified for some latent variable models.
In particular, in this work we focus on the probabilistic principal component analysis model to describe variation among treatments:
$U$ $=\epsilon_{U}$ (8)
$T$ $=BU+\epsilon_{T}$ (9)
where $B\in\mathbb{R}^{k\times m}$, $\epsilon_{U}\sim N_{m}(0,I_{m})$ and $\epsilon_{T}\sim N_{k}(0,\Sigma_{t|u})$ with $\Sigma_{t\mid u}$ being a diagonal matrix. This model implies the conditional confounder distribution is
@@
This leads to our main assumption, which will allow us to identify bounds on the causal effect.
@@
The treatments are generated by (8)–(9) and $U$ represents potential confounders in the sense that they are possible causes of $T$ and $Y$ and are not themselves caused by $T$ or $Y$. Additionally, we assume the following conditions are satisfied.
  1. 1.
@@
Given Assumptions 14, both $\|a^{\top}\Gamma\|_{2}^{2}$ and $\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|^{2}_{2}$ are identifiable and thus the bound on $\text{Bias}_{a,t_{1},t_{2}}$, given by $\|a^{\top}\Gamma\|_{2}^{2}\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|^{2}_{2}$, is also identifiable. Further, for any orthogonal matrix $\widetilde{\Gamma}$ satisfying $\widetilde{\Gamma}\widetilde{\Gamma}^{\top}=\Gamma\Gamma^{\top}$ and $\tilde{B}$, satisfying $\widetilde{B}\widetilde{B}^{\top}=BB^{\top}$ there exists an $m\times m$ orthonormal matrix, $R$, such that
$\text{Bias}_{a,t_{1},t_{2}}=a^{\top}\widetilde{\Gamma}R\widetilde{\Sigma}_{u|t}^{-1/2}\widetilde{\mu}_{u|\Delta t},$
where $\widetilde{\mu}_{u|\Delta t}$ and $\widetilde{\Sigma}_{u|t}^{-1/2}$ are the matrices implied by the particular value of $\tilde{B}$. The equivalence class of matrices $\widetilde{\Gamma}$ and $\widetilde{B}$ are identifiable, so the only unidentified component of the bias is $R$.
@@
Note that Assumption 4 states that $U$ are in fact confounders. The factor models reflect residual correlation in the treatments and outcomes, which we assume is driven entirely by unmeasured confounders. While this correlation could also be due to nonconfounding sources (e.g., predictors of only treatment or only outcome, or mediators), the bias bound in Theorem 1 still provides valid conservative bounds on the causal effects, as we will over-estimate the extent to which unmeasured confounding is biasing our estimates.
5 Joint Partial identification regions under additional constraints